OpenStudy (anonymous):

Short Answer. The length of the side of a rectangular desk is given by 3x-y cm. The length of the front is given by 3x+ y cm. What is an algebraic expression for the area of the top of the rectangular desk? Include correct units as part of your answer. Is the equation (3x-y)(3x+y)?

OpenStudy (anonymous):

That is the correct equation, 9x^2 - y^2 cm^2

OpenStudy (anonymous):

so you have to add the centimeters squared?

OpenStudy (anonymous):

yup. even for an equation you have to add that.

OpenStudy (anonymous):

(3x-y)(3x+y) F= 3x*3x = 9x^2 O=3x*y=3xy I=-y*3x= -3xy L=-y*y=-y^2 =9x^2 +3xy-3xy-y^2 =9x^2-y^2

OpenStudy (anonymous):

oh yeah =9x^2-y^2 cm^2